OpenStudy (anonymous):

6/2(1+2)=? first, (1+2)=(3) next 6/2=3 it looks like this,3(3) last, 3*(3)= 9 The answer is 9 does this look right?

OpenStudy (cggurumanjunath):

6/2(1+2) 6/(2*3) 6/6 1

OpenStudy (anonymous):

9 is right

OpenStudy (cggurumanjunath):

FOLLOW BODMAS RULE.

OpenStudy (anonymous):

you need to use BODMAS + left to right operator precedence here

OpenStudy (anonymous):

i got a 13/20 on it what i do wrong?

OpenStudy (cggurumanjunath):

THEN Q MUST BE (6/2)*(1+2)

OpenStudy (anonymous):

(6)/(2)*(1+2) Reduce the expression (6)/(2) by removing a factor of 2 from the numerator and denominator. 3*(1+2) Add 2 to 1 to get 3. 3*(3) Multiply 3 by each term inside the parentheses. 9

OpenStudy (texaschic101):

6/2 (1 + 2) 6/2 * 3 = 18/2 = 9

OpenStudy (anonymous):

(6/2)*(1+2) and 6/2(1+2) are same. thats the point i wana make :)

OpenStudy (texaschic101):

yes...I agree...they are the same

OpenStudy (anonymous):

so i did something wrong? im confused?

OpenStudy (anonymous):

your answer is correct

OpenStudy (anonymous):

funny my teacher gave me a 13 out of 20 on it

OpenStudy (anonymous):

ask ur teacher, dont hold back :)

OpenStudy (anonymous):

thanks 022

OpenStudy (anonymous):

u wlc 016 (:

OpenStudy (anonymous):

u too 101

OpenStudy (anonymous):

i have history no to do bla bla

OpenStudy (anonymous):

now*